silvernight269:

Before district play, the Unicorns had won 45% of their basketball games. During district play, they won six more games and lost two, to finish the season having won half their games. How many games did the Unicorns play in all?

silvernight269:

@TheSmartOne

TheSmartOne:

So let's say they played x games before the district, out of which they won w After the district, they've played a total of x + 8 games Before their winning ratio was 45% but after these 8 games, it became 50% So let's try to find x so then that we can find x+8 So at the end they ended up winning half their games So they played x games, once they won 6 more of them it became 50% (w + 6)/(x + 8) = 0.5 Remember 50% is the same thing as 0.5 And the other equation we have is w/x = 0.45

TheSmartOne:

Does this make sense so far

TheSmartOne:

So basically, you have two equations (w+6)/(x+8) = 0.5 and w/x = 0.45 You need to solve the system of equations for x. You can do this by multiplying by the denominator on both sides so that way you don't have to deal with fractions And then solve for one variable either through the elimination method, or the substitution method. Or the graphing method but that method is reserved for if you're already proficient at the first two and need to save time ;b